OpenStudy (anonymous):

\[{dy \over dx} = {\sqrt{x^2+y^2} - x \over y}\] Solve the DE by using a substitution

OpenStudy (experimentx):

perhaps, let y=vx

OpenStudy (anonymous):

This equation is homogenous, correct?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Divide by x in numerator and denominator and do as what @experimentX told :)

OpenStudy (anonymous):

divide by x?

OpenStudy (anonymous):

Yes \[{dy \over dx} = {\sqrt{1+(y/x)^2} - 1 \over (y/x)}\] Now proceed :)

OpenStudy (anonymous):

okay, I did that and then took y=vx so v=y/x and dy/dx = v + x dv/dx

OpenStudy (anonymous):

\[v + x{dv \over dx} = {\sqrt{1 + v^2} - 1 \over v}\] and got this

OpenStudy (anonymous):

Take v to RHS and then take LCM

OpenStudy (experimentx):

now use variable separation technique.

OpenStudy (anonymous):

I am getting log x = log (1- (y/x)) +log c

OpenStudy (anonymous):

\[{v \over \sqrt{1 + v^2} - 1 - v^2} dv = {dx \over x}\] is what I came up with

OpenStudy (anonymous):

@gmer , yes. take - sign out to write it as -(1+v^2) Then take t^2 = 1+v^2 and integrate

OpenStudy (experimentx):

I rhink 1 +v^2 = t would be better supposition, eliminates a step.

OpenStudy (anonymous):

@experimentX , no it rather simplifies the problem :D

OpenStudy (anonymous):

it --> t^2 = 1+ v^2

OpenStudy (anonymous):

but then how do you find dt if you use t^2 ?

OpenStudy (anonymous):

Simple 2 t dt = 2 v dv tdt = vdv

OpenStudy (anonymous):

Finally you get \[\int\limits_{}^{}{ tdt \over t -t^2 } = \int\limits_{}^{}{ dt \over 1 -t }\]

OpenStudy (anonymous):

@gmer , I hope you can complete the problem from here :D

OpenStudy (experimentx):

looks like your supposition was better than mine ... LOL

OpenStudy (experimentx):

gotta sleep good night!!

OpenStudy (anonymous):

Yes. It is 7 am. I better catch up with some sleep too. :D

OpenStudy (experimentx):

same here LOL:D

OpenStudy (anonymous):

Looks like many Indians have got addicted to OS :P

OpenStudy (anonymous):

Thanks guys, think that I got it

OpenStudy (anonymous):

Hey @shivam_bhalla , mind checking over my work? It got kinda weird at the end and I'm unsure of it